# Interpretable Spatio-Temporal Features Extraction based Industrial Process Modeling and Monitoring by Soft Sensor

### v1

@@ Abstract
Keywords:
Power plant, Soft sensor, Deep-learning, physics-informed modeling
@@ I Introduction
Process modeling and monitoring are the cornerstones of stable operation and security in large systems engineering. However, data offset, noise disturbance, low reliability of analyzers, and delays in long time-series data from most processes lead to a mismatch between models and the data, resulting in potential safety hazards and poor product quality. As a solution, soft sensors are used for process monitoring, quality prediction, and many other important applications [1].
@@
There are two main types of approaches to establish soft sensing models, namely, mechanism-based [2] and data-driven methods [3, 4]. Mechanism-based methods, also called first-principle models, use physicochemical governing equations and dynamic differential equations to describe industrial processes. When the mechanism of the process is well understood or sufficient knowledge about the process is available, mechanism-based methods can work effectively. However, the complexity of large-scale industrial systems often makes these preconditions difficult to meet, limiting the applicability of the first principles. In contrast, data-driven methods, particularly deep learning (DL), have achieved significant success in chemical, biochemical, and metallurgical processes, as they do not require extensive prior knowledge [5].
Due to the strong local spatio-temporal features of long time-series data in industrial processes, convolutional neural networks (CNNs) are widely applied in soft sensors for local feature learning [6]. For example, a variable correlation analysis-based CNN is proposed for topological feature extraction [7, 8], as demonstrated in the hydrocracking process and the debutanizer column. A dynamic CNN strategy is designed to learn hierarchical local nonlinear dynamic features for soft sensor modeling [9]. Furthermore, a multiscale attention-based CNN is proposed in [10] to extract multi-scale local features from complex multi-coupled process data. However, the inability of CNNs to capture global features and the potential far topological structure of variables limits the prediction accuracy of models. Hence, other deep learning models, such as long short-term memory (LSTM) [11], and transformers [12, 13] are applied in soft sensor. In [14], a novel multi-view spatial-temporal transformer network is proposed to learn complex spatial-temporal domain correlations and potential patterns from multiple views for accurate traffic-flow prediction. By incorporating an additional reconstruction constraint for the raw input data in each layer, a stacked enhanced auto-encoder is designed in [15] to learn features. However, the learned spatio-temporal features always assume spatio-temporal alignment, which ignores the time delays and coupling among variables, resulting in poor generalization.
@@
Another problem is that these black-box models lack physical interpretability. The captured spatio-temporal features of long time-series data are sparse, which cannot be explained by first principles, limiting their practicality. Although graph neural networks are considered a way to incorporate prior knowledge into networks, they are used primarily to identify physical correlations among variables [16, 17, 18]. Fortunately, a potential solution to this problem lies in physics-informed or guided neural networks, which are designed to enhance the transparency and explainability of deep learning models (PINNs) [19], ensuring that their outputs adhere to physical laws [20]. By integrating differential equations and initial conditions into the loss function, these models leverage the automatic differentiation capability of neural networks while embedding fundamental physical laws [21]. They have been successfully applied to lake temperature prediction [22], parameter estimation [23], and flexibility analysis [24]. While physics-informed modeling methods demonstrate advantages over traditional algorithms in efficiency, noise immunity, and accuracy, a core obstacle remains: apart from the loss function with physical laws, these deep learning methods are still black box models. The physical constraints in loss functions merely implicitly imply the presence of physical properties within the neural networks rather than providing explicit expressions of these relationships.
@@
Considering this gap, the primary objective of this study is to introduce an interpretable spatio-temporal feature extractor for soft sensors in industrial process modeling and monitoring. We propose a spatio-temporal consistent and interpretable model (STCIM) with an encoder-decoder structure and an interpretable latent space. This method unfolds in two distinct phases: in the first stage, a root layer is designed to extract the local features at different times, and an adaptive position encoding layer is embedded into the root layer to ensure the spatio-temporal consistency. Subsequently, the spatio-temporal features are utilized as input to a far topological alignment layer, computing the correlations among the features. The potential far topological coupling will be captured to calculate the latent space. In the second stage, the latent space is guided by the first principle and integrated into the loss function to predict the output. The latent space is explicitly assigned physical properties to maintain the interpretability of the model. Notably, this methodology holds potential for adaptation to other domains within industrial modeling or diagnostic applications. The efficacy of these concepts is validated through extensive experimentation using simulated datasets and real-world data from a 330MW power plant. In summary, the key contributions of this paper encompass:
@@
    A spatio-temporal consistent and interpretable model (STCIM) is proposed for industrial process modeling and monitoring. It efficiently extracts and aligns the spatio-temporal features of industrial processes and gives the latent variables physical meanings to improve the model’s interpretability.
@@
    The far topological spatio-temporal consistency extraction block is introduced to extract and align spatio-temporal features by using adaptive position encoding and attention modules.
@@
    The physics-informed interpretable latent space block is introduced to map the features into the state-space by giving explicit physical meanings to latent variables and integrating them in a discrete state-space loss function, enhancing the feature interpretability.
@@
    The efficacy of STCIM is underscored through its application experiments and ablation experiments in two simulated datasets and a real-world dataset, thus substantiating its effectiveness in practical settings.
The paper is organized as follows: Section II provides background about attention blocks, shortcut connections, and physics-informed loss functions. Section III describes the detailed information of the proposed STCIM including the far topological spatio-temporal consistency extraction block, the physics-informed interpretable latent space block, and the model structure. Section IV shows the experimental results, and Section V concludes the paper.
II PRELIMINARIES
II-A Attention Block
The attention block is widely used for large language models to measure the importance of each input word [25]. An attention function can be described as mapping a query and a set of key-value pairs to an output, which is shown in Figure 1a. The computation can be described as follows:
$Attention(Q,K,V)=SoftMax(\frac{QK^{T}}{\sqrt{d_{k}}})V$ (1)
where Q, K, and V are separate sets of queries, keys, and values. $\frac{1}{\sqrt{d_{k}}}$ is the scaling factor. In particular, when we use the self-attention block, Q, K, and V are the same. When the input is time-series data, it can extract intrinsic temporal correlations on the timeline.
(a) Attention Block.
(b) Residual Block.
Figure 1: The basic attention (left) and residual (right) blocks [25].
II-B Shortcut connections
Difficulty introduced by the degeneracy phenomenon is always encountered in training neural networks with large depth. One of the possible solutions is the shortcut connection (residual connection), which has had great success in many applications [26]. In the aspect of information theory, the introduced shortcut connections in residual neural networks enable input data to flow directly to the next layer without information loss. It has been proven that residual neural networks can improve the performance of networks with an increase in depth. The basic residual block used in Transformer is shown in Figure 1b, and can be expressed as
$y^{l}=F(y^{l-1},{W^{l},b^{l}})+y^{l-1}$ (2)
where $y^{l}$ and $y^{l-1}$ are the output of layer $l$ and $l-1$.$W^{l},b^{l}$ are the weights and bias of layer $l$. $F(\cdot)$ is the residual function that contains the linear transformations and the Gaussian error linear unit (GELU) activation functions [27].
II-C Physics-informed loss function
The physics-informed loss function is designed based on physical systems and available datasets. Assuming that the physical system can be represented as follows:
$\begin{split}&\text{Physical Model:}y_{t,x}=-f(u;\theta)(x),\\ &\text{Initial Conditions:}\Psi(u;\theta)(x,t_{0})=0,\\ &\text{Boundary Conditions:}\Gamma(u;\theta)(x,t)=0.\end{split}$ (3)
where $y_{t,x}$ is the output of the model, $f$ is the nonlinear function, $x\in\Omega$ is the input of model, $\Omega\in\mathbb{R}^{d}$, $u\in\mathbb{R}^{m}$ are the state variables of the system, $\theta$ is the parameters of model, and $t\in[t_{0},T]$ is the time domain (all variables and parameters are vectors). For dynamic systems, initialization and boundary conditions are defined separately as $\Psi(u;\theta)$ and $\Gamma(u;\theta)$.
Considering the above physical system and assuming that we have a dataset $\mathcal{D}$ containing all state variables $u(x)$, input variables $x$, time $t$, and output variables $y$, the physics-informed loss function is defined as follows:
$\begin{split}Loss=L(y,\hat{y})+\lambda R(\theta)+\gamma R_{Phy}(x,y)\end{split}$ (4)
where $\hat{y}$ is the actual data, $L$ is the Mean Square Error (MSE), $R(\theta)$ is the parametric regularization of the model (i.e. $L1$ and $L2$ norm regularization), $R_{Phy}(x,y)$ is the physical constraint loss which consists of physical models, initial conditions, and boundary conditions (i.e $R_{Phy}(x,y)=y_{t,x}+f(u;\theta)(x)$). $\lambda$ and $\gamma$ are separately the hyper-parameters of the parametric regularization and physical constraint. According to Eq. (4), the physical prior knowledge in Eq. (3) is embedded in the loss function to constrain neural networks.
@@ III Spatio-Temporal Consistent and Interpretable Modeling
Modern multi-coupled industrial processes are composed of multiple interconnected production units that have complex material and energy transfers and intricate structures. Hence, time delay and coupling among variables in industrial processes are the main obstacles for modeling and monitoring. In this section, we introduce the proposed far topological spatio-temporal consistency extraction block and the physics-informed interpretable latent space block for the interpretable feature extraction in soft sensors.
III-A Far Topological Spatio-Temporal Consistency Extraction Block
The far topological spatio-temporal consistency extraction block can be divided into two layers: the spatio-temporal feature extraction layer and the far topological alignment layer.
III-A1 Spatio-Temporal Feature Extraction Layer
Figure 2 describes the spatio-temporal feature extraction layer. The input data, respectively, pass through a temporal feature extraction layer and an adaptive position encoding layer to capture the local features from different times and locations. The massive local spatio-temporal features ensure the abundance of far topological variables.
@@ III-A1 Spatio-Temporal Feature Extraction Layer
where $T$ is the sampling time. When different convolution kernels are applied in $X(t)$, the local features of adjacent attributes and times are extracted, reflecting the dynamics of the industrial process. However, these features come from different attributes at different times, lacking the corresponding spatial features. To overcome the problem, inspired by [28], an adaptive position encoding method is given by
$Position_{i}=\frac{Variable_{i}}{\sqrt{N}}$ (6)
@@
where $Variable_{i}$ is the trainable position parameter of $i_{th}$ attributes. $N$ is the number of features that is a scaling factor. $Position_{i}$ is the position encoding of the $i_{th}$ attributes. Due to the unknown time delay and coupling among variables in industrial processes, fixed position encoding is not appropriate, for example, ’sin’ and ’cos’ encoding. The trainable position parameter $Variable_{i}$ in this layer can adaptively ensure spatio-temporal consistency of captured local features.
Figure 2: Spatio-temporal consistent block.
III-A2 Far Topological Alignment Layer
@@
During the industrial process, variables are related not only to adjacent nodes but also to variables in far topological structures. Aligning the topological correlation among spatio-temporal local features is vital for industrial process modeling and monitoring. In this vein, we propose a far topological alignment layer, which is shown in Figure 3. This layer is coupled with the spatio-temporal feature extraction layer. Local spatio-temporal features are sparsely mapped and aligned by multi-head attention modules to capture the topological correlation in industrial processes. The calculated attention weights can be considered as the dependencies among far topological features and captured by models. Subject to potential far topological variable correlations, the adaptive position encoding at the spatio-temporal feature extraction layer is realigned with new features, which will again improve the modeling and monitoring accuracy.
@@
Assume that the features captured by the spatio-temporal feature extraction layer are $V\in R^{H\times W\times K}$. Based on the attention weights in Eq. (1), the dependencies among far topological features can be calculated by
$Weight(V,V)=SoftMax(\frac{VV^{T}}{\sqrt{d_{k}}})$ (7)
@@
where $Q=K=V$. The calculated weights determine the intrinsic correlation among the input features, helping the networks capture the potential far topological structure among all the features. By multiplying $Weight(V,V)$ with $V$, the importance determined by the long-term topological correlation can be used in feature extraction.
Figure 3: Far topological structure capture block.
III-B physics-informed Interpretable Latent Space Block
@@
To limit the captured features in first principles and map these sparse features into an interpretable state-space, this subsection introduces the physics-informed interpretable latent space block by giving physical meanings to the latent space $z$ of STCIM. We can directly relate the inputs, latent variables, and outputs by the physics-informed discrete state-space loss function and boundary conditions. Here, we need to highlight that the latent space $z$ is an intermediate process parameter instead of the input of STCIM. It can be unobserved or hard to measure.
@@
$u(k)=\Phi(u(k-1),\theta_{1})+G(x(k),\theta_{2})$ (8)
$y(k)=H(u(k),\theta_{3})+J(x(k),\theta_{4})$ (9)
where $x$ is the input, $u$ is the unobserved state variables, and $y$ is the output. $\Phi(\cdot),G(\cdot),H(\cdot),J(\cdot)$ are separately the nonlinear operators. $\theta$ is the parameter. All variables are vectors. Now, we can map the discrete-time state-space equation into STCIM. Assuming that the latent space $z$ of STCIM is the state variables at time $k-1$ and $k$ ($z_{1}=u(k)$, $z_{2}=u(k-1)$ and $z=[z_{1},z_{2}]$) and the input and output of the STCIM are $x$ and $\hat{y}$, respectively. Then, the discrete state-space loss function of STCIM can be given by
$\begin{split}Loss=L(y,\hat{y})+\lambda R(\theta)+\gamma_{1}(\hat{y}-H(z_{1},\theta_{3})-\\ J(u,\theta_{4})+\gamma_{2}(z_{1}-\Phi(z_{2},\theta_{1})-G(x,\theta_{2}))\end{split}$ (10)
where $\gamma$ is the weight coefficient. When the $\gamma=[\gamma_{1},\gamma_{2}]\to\infty$, the STCIM completely trusts the state-space model. In contrast, when the $\gamma=[\gamma_{1},\gamma_{2}]\to 0$, the STCIM is a data-driven model without physical prior knowledge. By balancing the tendencies towards data-driven and physics-driven models, STCIM can achieve the trade-off in the loss function. According to Eq. (10), the discrete state-space loss function consists of four sub-losses: 1) the MSE loss $L(y,\hat{y})$, 2) the parametric regularization of the model $R(\theta)$, 3) the output equations loss $\hat{y}-H(z_{1},\theta_{3})-J(x,\theta_{4}$, 4) the state equation loss $z_{1}-\Phi(z_{2},\theta_{1})-G(x,\theta_{2})$. Although the state variables are not measurable, the neural network will output the actual state variables because the latent space is constrained at both sub-losses 3 and 4, which provides physical properties.
III-C Model Structure
Figure 4 shows the structure of the spatio-temporal consistent and interpretable model. It can be recognized as an encoder-decoder structure, which is generally considered as a feature extractor and a function fitter when applied in many aspects. The idea of the model can be summarized as compressing the spatio-temporal consistent features with high dimensions into an interpretable latent space, which normally has fewer dimensions, and then the interpretable latent variables are utilized to fit the potential mathematical equations.
As aforementioned, the far topological spatio-temporal consistency extraction block is used as the encoder to capture and align the far topological spatio-temporal features among variables. The decoder is composed of MLPs, which is enough for discrete function fitting. The latent variables and the outputs are endowed with physical meanings by using a physics-informed interpretable latent space block. The size of the latent space is determined by the form of discrete state-space equations, e.g., if there are two state variables in a discrete model, then the number of latent variables for STCIM is four because it contains the state variables at time $k$ and $k-1$.
Figure 4: Spatio-temporal consistent and interpretable model.

### v2

# Spatio-Temporal Consistent Soft Sensor Modeling and Monitoring of Thermal Power Plants based on Physical Knowledge

@@ I Introduction
To reduce carbon emissions and mitigate the greenhouse effect, the modeling and monitoring of thermal power plants become the cornerstones of stable operation and security in the energy system. However, data offset, noise disturbance, low reliability of analyzers, and delays in long time-series data from most processes lead to a mismatch between models and the data, resulting in potential safety hazards and poor product quality. As a solution, soft sensors are used for process monitoring, quality prediction, and many other important applications in thermal power plants [1].
@@
There are two main types of approaches to establish soft sensing models, namely, mechanism-based methods [2] and data-driven methods [3, 4]. Mechanism-based methods, also called first-principle models, use physicochemical governing equations and dynamic differential equations to describe industrial processes [5, 6]. It can model or monitor thermal power plants in an interpretable and reliable way. When we have sufficient knowledge about thermal power plants and the process, mechanism-based methods can work effectively, improving practitioners’ trust in the models. However, the complexity of thermal power plants often makes these preconditions difficult to meet, limiting the applicability of the first principles. In contrast, data-driven methods, particularly deep learning (DL), have achieved significant success in chemical, biochemical, and metallurgical processes, since they do not require extensive prior knowledge [7, 8, 9].
In recent research, deep learning methods have been increasingly applied to large-scale industrial processes, with a particular emphasis on spatio-temporal feature extraction to capture both spatial dependencies and temporal dynamics among variables [10]. These approaches build on foundational techniques such as multiscale information fusion [11] and dynamic feature extraction [12], but prioritize spatio-temporal integration to address complex interactions. Some efforts focused on local spatio-temporal features [13], incorporating mechanisms like BiGRU with attention for time-related patterns [14] and dynamic convolutional neural networks with multiscale attention for hierarchical local nonlinear dynamics [15]. However, the far topological structures in industrial variables often limit local methods, prompting advancements in broader spatio-temporal modeling [16]. For instance, variable correlation analysis-based convolutional neural networks have been proposed to extract far topological spatio-temporal features in processes like hydrocracking and debutanizer columns [17]. Similarly, spatio-temporal attention combined with long short-term memory networks identifies long-term variable correlations in hydrocracking [18], while graph neural networks process relational structures for enhanced spatio-temporal feature extraction [19]. However, many learned broader spatio-temporal features assume alignment across space and time, overlooking delays and couplings among variables, which can lead to reduced generalization in real-world applications. This problem is particularly evident in thermal power plants and needs to be solved.
@@
Another problem is that black-box models lack physical interpretability, critical for safety in thermal power plants. Sparse spatio-temporal features in long time-series data are not explainable by first principles, limiting their utility. While graph neural networks incorporate prior knowledge to identify physical correlations [20, 10], they lack explicit physical grounding. Physics-informed neural networks (PINNs) address this by embedding physical laws into loss functions via differential equations and initial conditions, leveraging automatic differentiation for transparency [21, 22, 23]. Successful applications include lake temperature prediction [24], parameter estimation [25], and flexibility analysis [26]. Despite advantages in efficiency, noise immunity, and accuracy, PINNs implicitly encode physical properties, lacking explicit expressions of these relationships.
@@
Considering this gap, the primary objective of this study is to introduce an explicitly interpretable spatio-temporally consistent soft sensor model in thermal power plant modeling and monitoring. We propose a spatio-temporal consistent and interpretable model (STCIM) with a spatio-temporal consistent feature extraction block and an interpretable latent space. This method unfolds in two distinct phases: in the first stage, a root layer is designed to extract the local features at different times, and a trainable position encoding layer is embedded into the root layer to ensure the local spatio-temporal consistency. Subsequently, the local consistent features are utilized as the input to a far topological alignment layer, computing the correlations among the features by attention mechanism. During the process, the position encoding layer can adaptively update itself to ensure the consistency of far topological spatio-temporal features. In the second stage, the spatio-temporal consistent features are projected into the latent space. The latent space is guided by the first principle, explicitly assigning physical meanings to maintain the interpretability of the model. In training process, the latent space is integrated into the loss function to predict the state variables and the output. Notably, this methodology holds potential for adaptation to other domains within industrial modeling or diagnostic applications. The efficacy of these concepts is validated through extensive experimentation using simulated datasets and real-world data from a 330MW thermal power plant. In summary, the key contributions of this paper encompass:
@@
    A spatio-temporal consistent and interpretable model (STCIM) is proposed for thermal power plant modeling and monitoring. It efficiently extracts and aligns the spatio-temporal features and gives the latent variables physical meanings to improve the model’s interpretability.
@@
    The spatio-temporal consistent feature extraction block is introduced to extract and align local and global spatio-temporal features by using trainable position encoding and attention modules.
@@
    The physics-informed interpretable latent space loss function is introduced to map the latent space into the state variables by giving explicit physical meanings to latent space and integrating them in a discrete state-space loss function. The discrete state-space loss function is a part of the physics-informed interpretable latent space loss function.
@@
    The efficacy of STCIM is underscored through its application experiments and ablation experiments in two simulated datasets and two real thermal power plant datasets, thus substantiating its effectiveness in practical settings.
The paper is organized as follows: Section II provides background about the thermal power plant, basic blocks, and physics-informed loss functions. Section III describes the detailed information of the proposed STCIM, including the model structure, the spatio-temporal consistent feature extraction block, and physics-informed interpretable latent space loss function. Section IV shows the experimental results, and Section V concludes the paper.
II Brief Introduction of Thermal Power Plants
Figure 1: 330MW thermal power plant.
Thermal power plant is a multifaceted and complex industrial process, generating electricity by converting heat energy from fossil fuels into electrical energy [27]. It contains the chemical reactions, energy conversion, mass transfer which is shown in Fig. 1. The process begins with fuel combustion in a boiler, producing high-pressure steam that drives a turbine connected to a generator, which produces electricity. After passing through the turbine, the steam is cooled in a condenser and recycled, while the exhaust gases are treated and released through a chimney.
The mathematical description of the utilized 330MW thermal power plant [28, 29] can be given by
where $u_{B}$ is the fuel signal ($kg/s$), $u_{T}$ is the opening of turbine ($\%$), $r_{B}$ is the actual fuel ($kg/s$), $p_{d}$ is the pressure of steam (MPa) and $N_{E}$ is the output power (MW). $\dot{z_{1}}=\frac{dr_{B}}{dt}$, $\dot{z_{2}}=\frac{dp_{d}}{dt}$ and $\dot{y}=\frac{dN_{E}}{dt}$ ($z_{1},z_{2}$ are the state variables). It is derived from the first principles of the thermal power plant.
@@ III Spatio-Temporal Consistent and Interpretable Modeling
Modern multi-coupled industrial processes involve interconnected production units with complex material and energy transfers, posing challenges in modeling and monitoring due to time delays, variable coupling, and interpretability. We developed a trainable position coding to align spatio-temporal consistency during local and far topological feature extraction. Additionally, we designed a physics-informed interpretable latent space loss function by assigning physical meanings to latent variables and incorporating them into the loss function. This section details the STCIM, focusing on the spatio-temporal consistent feature extraction block and the physics-informed interpretable latent space loss function.
Figure 2: Overview of the STCIM framework. (a) The flow chart of STCIM. (b) The local spatio-temporal feature extraction block. (c) The far topological alignment block. (d) The physics-informed interpretable latent space loss function
III-A Model Structure
Fig. 2 (a) shows the flow chart of the proposed STCIM. It can be recognized as an encoder-decoder structure, in which the spatio-temporal consistent feature extraction block (encoder) is divided into two sub-blocks: the local spatio-temporal consistent feature extraction block and the far topological alignment block, and the decoder is a fully connected network.
In the local spatio-temporal consistent feature extraction block, the temporal features are extracted by the kernels with different size. The trainable position layer is added in parallel to temporal feature extraction as the position embedding to ensure the consistency of local spatio-temporal features. Then the local features are utilized to calculate the far topological correlations based on the attention mechanism, capturing the coupling among features in the far topological alignment block. During the process, the position encoding can adaptively update itself based on the new captured coupling features to ensure the consistency of far topological spatio-temporal features.
The latent space is another important node, which is utilized to explicitly add physical meanings to models. The spatio-temporal consistent features are projected into the latent space and constrained by the first principles by adding a discrete state-space loss in the physic-informed interpretable latent space loss function. The size of the latent space is determined by the form of discrete state-space equations, e.g., if there are two state variables in a discrete model, then the number of latent variables for STCIM is four because it contains the state variables at time $k$ and k1k-1. The decoder is composed of MLPs, which is enough for discrete function fitting.
In general, the idea of the model can be summarized as compressing the spatio-temporal consistent features with high dimensions into an interpretable latent space, which normally has fewer dimensions, and then the interpretable latent variables are utilized to fit the potential mathematical equations, for example, the state-space equations.
III-B Spatio-Temporal Consistent Feature Extraction Block
The spatio-temporal consistent feature extraction block mainly focuses on the consistency of space and time in features by designing a trainable position encoding layer. It contains the local spatio-temporal consistent feature extraction block and the far topological alignment block. In the two blocks, during the training process, the position encoding layer can adaptively update itself to ensure the spatio-temporal consistency in any features.
III-B1 Local Spatio-Temporal Feature Extraction Block
Fig. 2 (b) describes the local spatio-temporal feature extraction block. The input data, respectively, pass through a temporal feature extraction layer and a trainable position encoding layer to capture the local features from different times and locations. The massive local spatio-temporal features ensure the abundance of far topological variables.
@@ III-B1 Local Spatio-Temporal Feature Extraction Block
where $T$ is the sampling time. When different convolution kernels are applied in $X(t)$, the local features of adjacent attributes and times $S_{i,j}$ are extracted, reflecting the dynamics of the industrial process, $S_{i,j}=\sum_{m=1}^{m}\sum_{n=1}^{n}K_{m,n}X(t)_{i+m,j+n}+b$.
However, this kind of one-shot input is disorder, ignoring the potential position information in attributes, which is opposite to the clear production procedures of the industrial process. Therefore, we designed a trainable position encoding layer, providing the corresponding spatial features to ensure the consistency of spatio-temporal features. Inspired by [30], the trainable position encoding layer is given by
$Position_{j}=\frac{Variable_{j}}{\sqrt{N}}$ (3)
@@
where $Variable_{j}$ is the trainable position parameter of $j_{th}$ attributes. $N$ is the number of features that is a scaling factor. $Position_{j}$ is the position encoding of the $j_{th}$ attributes. Due to the unknown time delay and coupling among variables in industrial processes, fixed position encoding is not appropriate, for example, ’sin’ and ’cos’ encoding. The trainable position parameter $Variable_{j}$ in this layer can adaptively ensure spatio-temporal consistency of captured local features.The final local spatio-temporal consistent features can be summarized as
$\hat{S_{j}}=S_{j}+Position_{j}$ (4)
where $\hat{S_{j}}$ is the final feature of $j_{th}$ attributes.
III-B2 Far Topological Alignment Block
@@
During the industrial process, variables are related not only to adjacent nodes but also to variables in far topological structures. Aligning the far topological correlation among local features and making sure the spatio-temporal consistency are vital for the modeling and monitoring of thermal power plant. In this vein, we propose a far topological alignment block, which is shown in Fig. 2 (c).
The local spatio-temporal consistent features are first sparsely mapped by the multi-head linear transformations to highlight different aspects of the features. Then several self-attention modules are utilized to capture the far topological structure in features by calculating the attention weights of all features. The calculated weights can be considered as the dependencies among far topological features and captured by models. Subsequently, the captured topological and spatio-temporal features pass through a fully connected module with GeLU to provide nonlinear properties. In the forward process, there are two shortcut connections to ensure the flow of information. During the far topological variable correlation extraction process, the adaptive position encoding at the local spatio-temporal feature extraction layer is realigned with new features, which will again ensure the consistency of the far topological spatio-temporal feature and improve the modeling and monitoring accuracy.
@@
Assume that the features captured by the local spatio-temporal feature extraction block are $\hat{S}\in R^{H\times W\times K}$. Based on the attention weights, the dependencies among far topological features can be calculated by
$Weight(\hat{S},\hat{S})=SoftMax(\frac{\hat{S}\hat{S}^{T}}{\sqrt{d_{k}}})$ (5)
$S_{new}=Weight(\hat{S},\hat{S})\hat{S}$ (6)
@@
where $Q=K=V=\hat{S}$. The calculated weights determine the intrinsic correlation among the input features, helping the networks capture the potential far topological structure among all the features. By multiplying Weight(S^,S^)Weight($\hat$$\hat{S}$) with $\hat{S}$, the importance determined by the long-term topological correlation can be used in feature extraction.
III-C Physics-informed Interpretable Latent Space Loss Function
@@
To map these sparse spatio-temporal consistent features into an interpretable latent space and limit the space in first principles, this subsection introduces the physics-informed interpretable latent space loss function by giving physical meanings to the latent space $z$ of STCIM and adding them to the discrete state-space loss, which is shown in Fig. 2 (d). We can directly relate the inputs, latent variables, and outputs by the discrete state-space loss function and boundary conditions. Here, we need to highlight that the latent space $z$ is an intermediate process parameter instead of the input of STCIM. It can be unobserved or hard to measure.
@@
$u(t)=\Phi(u(t-1),\theta_{1})+G(x(t),\theta_{2})$ (7)
$y(t)=H(u(t),\theta_{3})+J(x(t),\theta_{4})$ (8)
where $x$ is the input, $u$ is the unobserved state variables, and $y$ is the output. $\Phi(\cdot),G(\cdot),H(\cdot),J(\cdot)$ are separately the nonlinear operators. $\theta$ is the parameter. All variables are vectors.
Now, we can map the discrete-time state-space equation into STCIM by defining the latent variables with state variables in Eq. (7). Assuming that the latent space $z$ of STCIM is the state variables at time t1t-1 and $t$ ($z_{1}=u(t)$, $z_{2}=u(t-1)$ and $z=[z_{1},z_{2}]$) and the input and output of the STCIM are $x$ and $\hat{y}$, respectively. Then, the discrete state-space loss function can be given by
$\begin{split}Loss_{state}=\gamma_{1}(\hat{y}-H(z_{1},\theta_{3})-J(x,\theta_{4}))+\\ \gamma_{2}(z_{1}-\Phi(z_{2},\theta_{1})-G(x,\theta_{2}))\end{split}$ (9)
where $\gamma$ is the weight coefficient. The discrete state-space loss function is used to constrain the latent variables to follow the state-space equations, i.e. the underlying laws of physics. By adding it to the MSE loss, we get the physics-informed interpretable latent space loss function, given by
$\begin{split}Loss=L(y,\hat{y})+\lambda R(\theta)+Loss_{state}\end{split}$ (10)
When $\gamma=[\gamma_{1},\gamma_{2}]\to\infty$, the STCIM completely trusts the state-space model. In contrast, when $\gamma=[\gamma_{1},\gamma_{2}]\to 0$, the STCIM is a data-driven model without physical prior knowledge. Assuming that the devices follow a unified non-linear model across different operating conditions, the optimal $\gamma$ is certain and determined by whether this model can be trusted,i.e., how simplified the state-space model is. By balancing the tendencies towards data-driven and physics-driven models, STCIM can achieve the trade-off in the loss function. According to Eq. (10), the physics-informed interpretable latent space loss function consists of four sub-losses: 1) the MSE loss L(y,y^)L($\hat{y}$), 2) the parametric regularization of the model $R(\theta)$, 3) the output equation loss $\hat{y}-H(z_{1},\theta_{3})-J(x,\theta_{4}$, 4) the state equation loss $z_{1}-\Phi(z_{2},\theta_{1})-G(x,\theta_{2})$. Although the state variables are not measurable, the neural network will output the actual state variables due to the latent space being constrained at both sub-losses 3 and 4, which provides physical properties.